A gardener has 67 feet of fencing to be used to enclose a rectangular garden that has a border 2 feet wide surrounding it​. If the length of the garden is to be twice the width, what will be the dimensions of the garden.
Answer by ankor@dixie-net.com(22740)

A gardener has 67 feet of fencing to be used to enclose a rectangular garden that has a border 2 feet wide surrounding it​.
If the length of the garden is to be twice the width, what will be the dimensions of the garden.
:
let w = the width of the garden
then
2w = the length
:
The dimension of the fenced area.
(w+4) by (2w+4)
the perimeter
2(w+4) + 2(2w+4) = 67
2w + 8 + 4w + 8 = 67
6w + 16 = 67
6w = 67 - 16
6w = 51
w = 51/6
w = 8.5 ft is the width of the garden
and
2(8.5) = 17 ft is the length